Please Help!
25= √5^(2x+6)
explain the strategy used in solving this equation the way you did

Answer by rothauserc(4718)

first note that sqrt(5) = 5^1/2
with that we can rewrite the equation
25 = (5^1/2)^(2x+6)
multiply the exponents on the right side of = sign
25 = 5^(x+3)
and x = -1 since 5^2 = 25